Gold (Au) is in the same group on the periodic tab
Answer #1

given,

density, d = 19.3 g/cm^3

atomic mass, m = 196.67 g/mol

No.of gold atoms per unit volume, n = (density/atomic mass)* Avogadro number

= (19.3/196.67) * 6.03*10^{23}

= 5.9174 *10^{22} electrons/cm^3

= 5.9174*10^{26} electrons/m^3

Relaxation time is

\tau =\frac{m\rho}{ne^2}=\frac{9.1*10^{-31}*4.1*10^7}{5.9174*10^{28}*(1.6*10^{-19})^2}

= 2.463*10^{-14} s

The electron mobility is

  \mu =\frac{e\tau}{m} = \frac{1.6*10^{-19}*2.463*10^{-14}}{9.1*10^{-31}}=4.3305*10^{-3}\:m^2/V-s